OpenStudy (thedj4jc):

If the quadratic formula is used to find the solution set of 3x2 + 4x - 2 = 0, what are the solutions? Answer options in replies

OpenStudy (thedj4jc):

A. \[\left\{ \frac{ -2\pm \sqrt{10} }{ 3 } \right\}\] B. \[\left\{ \frac{ -2\pm \sqrt{2} }{ 3 } \right\}\] C. \[\left\{ \frac{ 2\pm \sqrt{10} }{ 3 } \right\}\]